### v1

@@ 6 Conclusion
In this work, we investigate the stability of quantum systems by analyzing the dynamics of CTQWs on various network topologies under different decoherence mechanisms. We define stability in terms of the preservation of quantum properties over time and examine how this preservation is influenced by the interplay between network structure and decoherence model. To analyze this, we consider both simple network structures, including star, cycle, and complete graphs, and complex topologies such as Erdős–Rényi, small-world, and scale-free networks.
@@
We evaluate the dynamics of CTQWs on these networks to examine the persistence of quantum properties, both in the absence of noise and under three decoherence models namely, Haken-Strobl, intrinsic decoherence and QSW. This systematic analysis allows us to explore how network structure and decoherence jointly influence CTQW stability. To quantify the stability, first we compute the node probability distribution of the quantum walker and the $\ell_{1}$-norm of coherence. In addition to these primary measures, we also use fidelity, quantum–classical distance, and von Neumann entropy measures for providing a comprehensive characterization of the preservation of quantum properties.
@@
Based on these metrics, we find that stability varies significantly across decoherence models, for example intrinsic decoherence preserves coherence most effectively, followed by Haken–Strobl noise, while QSW result in the most rapid coherence loss. Our results reveal that coherence preservation depends not only on the type of decoherence but also on the underlying topology and the specific node at which the quantum walker is initialized.
@@
Delving deeper, we observe that homogeneous networks such as the cycle exhibit high coherence in the absence of noise but are highly vulnerable under decoherence. In contrast, heterogeneous topologies like the star network demonstrate greater robustness under Haken–Strobl and intrinsic decoherence, as quantified by the $\ell_{1}$-norm of coherence. In the case of QSW, star networks initialized at peripheral nodes retain coherence longer and delay the onset of uniform probability distribution compared to central hub initialization. Interestingly, complete and star networks initialized at central nodes exhibit similar probability distributions and similar stability measures under decoherence, reflecting the impact of local connectivity due to comparable node degrees.
@@
For complex networks, when coherence preservation is quantified using $\ell_{1}$-norm of coherence indicates that scale-free networks are the most stable under Haken–Strobl but exhibit slightly faster coherence decay than Erdős–Rényi networks under QSW dynamics. When stability is assessed using fidelity and von Neumann entropy, we observe distinct temporal behaviors across Erdős–Rényi, small-world, and scale-free networks. These results indicate that the increased heterogeneity of scale-free networks contributes to their enhanced stability. However, when using the quantum–classical distance as a stability measure, particularly under noiseless and intrinsic decoherence case, all network types yield nearly identical trajectories, making it difficult to distinguish relative stability. The interpretation of stability is inherently application-dependent. For example, high entropy may impair the efficiency of quantum search algorithms [50], while a large quantum–classical distance, indicating strong deviation from classical dynamics can benefit the design of optimized chiral quantum walks [51].
In general, heterogeneous networks exhibit greater coherence stability than homogeneous networks with less connectivity. Moreover, the stability of such networks is strongly influenced by the initialization of the quantum walker at high-degree nodes, which tends to promote localization and coherence retention under noise but reduces coherence in noiseless conditions. This observation highlights a fundamental trade-off between localization and coherence. The implications of this trade-off are particularly relevant for quantum information processing: coherence can serve as a resource for quantum search algorithms based on quantum walks [52, 53], whereas high fidelity often indicative of localization can be exploited for robust quantum memories [54].
Based on our analysis, there is significant scope for future research on quantum coherence stability in noisy environments by investigating additional structural properties of networks and considering the dynamics of multiple continuous-time quantum walkers (multi-CTQWs).
@@ Acknowledgments
ALJ acknowledges the University Grants Commission (UGC), India, for financial support. CM acknowledges support from the Anusandhan National Research Foundation (ANRF) India (Grants Numbers SRG/2023/001846 and EEQ/2023/001080).
Appendix A CTQW on Triangle Graph
Figure 8: Triangle graph with the initial localization of the CTQW at node 00.
For the triangle graph, we consider that the CTQW is localized at node 00, as illustrated in Fig. 8. The system’s Hamiltonian is given by the graph Laplacian:
$H=\begin{bmatrix}2&-1&-1\\ -1&2&-1\\ -1&-1&2\end{bmatrix}$ (12)
In the noiseless case, the evolution of the density matrix is governed by the unitary transformation according to :
$\rho(t)=e^{-iHt}\rho_{0}e^{iHt}$ (13)
For the triangle graph, this results in the following density matrix, derived analytically:
$\rho_{t}$ $=$ (14)
$\frac{1}{9}\begin{bmatrix}5+4\cos{3t}&-1+2e^{-3it}-e^{3it}&-1+2e^{-3it}-e^{3it}\\ -1+2e^{3it}-e^{-3it}&2-2\cos{3t}&2-2\cos{3t}\\ -1+2e^{3it}-e^{-3it}&2-2\cos{3t}&2-2\cos{3t}\end{bmatrix}$
From Figs. 9(a),(e), we observe that the node occupation probabilities obtained from the theoretical denoted as a dashed line, match exactly with the occupation probability extracted from the numerical density matrices, confirming the consistency between the analytical and simulation results.
As we know, the long-time behavior of node occupation probabilities serves as a reliable indicator of quantum stability. From Fig. 9, we observe that in the absence of decoherence, CTQWs on the triangle graph exhibits persistent oscillatory dynamics without reaching a steady probability distribution, reflecting sustained quantum behavior. In contrast, the introduction of decoherence drives the system toward a steady probability distribution, with the rate and nature of convergence depending critically on both the decoherence model and the network topology.
As shown in Fig. 9(b),(f), the presence of Haken–Strobl noise leads to damped oscillations in the node occupation probabilities, which gradually converge to a uniform distribution. This behavior closely resembles that of the QSW model, Fig. 9(d),(h), however, in the QSW, convergence is faster and smoother, with oscillations decaying over a shorter timescale. In contrast, under intrinsic decoherence, Fig. 9(c),(g), the occupation probabilities do not evolve toward uniformity. Instead, they stabilize at two distinct values, highlighting a fundamental difference in the asymptotic behavior across these decoherence models.
These observations indicate that, for the triangle graph, intrinsic decoherence tends to preserve quantum coherence more effectively than both the Haken–Strobl and quantum stochastic walk models. This underscores the critical role of the decoherence mechanism in shaping the long-term quantum dynamics of the system.
Figure 9: Occupation probabilities for quantum walker in a triangle graph. The (a)-(d) shows the time evolution of the node occupation probabilities, with numerical results matching the analytically derived density matrix solutions (gray dashed lines). Nodes 2 and 3 exhibit identical probabilities, leading to overlapping curves. (e)-(h) presents the long-time occupation probabilities, using the same color scheme as in the top panels.
Appendix B CTQWs on Network Topologies ranging from Homogeneous to Heterogeneous structures
To further investigate which network structures are most effective in preserving quantum properties and supporting the stability of quantum systems, we analyze CTQWs on several network topologies, ranging from homogeneous structures such as the cycle graph to highly heterogeneous ones such as the star network. The objective is to identify topologies that offer the greatest stability in various decoherence models.
To investigate this, we calculate the quantum-classical distance, fidelity with the initial state, and von Neumann entropy for the cycle, star, complete, star, Erdős–Rényi, small-world and scale-free network topologies. Results for networks of size $N=10$ are presented in Fig. 10(a)–(h) and Fig. 11(a)–(c), while results for network size $N=50$ are shown in Fig. 12(a)–(h) and Fig. 13(a)–(c).
These figures show that simulations with larger network sizes reveal a clearer distinction in the dynamical behavior across different topologies. From these figures, we observe that for all decoherence models, except the QSW, the complete graph (red curves) and the star network (orange curves) with hub node initialization consistently demonstrate the highest stability. This is evidenced by their highest quantum-classical distance and fidelity with respect to the initial state, along with the lowest von Neumann entropy. These findings are consistent with earlier results based on the $\ell_{1}$-norm of coherence.
In contrast, the star network with peripheral node initialization tends to classicalize more rapidly, as indicated by its lowest quantum-classical distance, suggesting reduced stability. However this same configuration displays a higher fidelity with the initial state and lower entropy compared to other topologies. This highlights that different stability metrics capture distinct aspects of the CTQW’s stability in different topologies and are relevant in different contexts, as elaborated in the conclusion (Sec. 6). Additionally, under the QSW decoherence, the decay of $\ell_{1}$-norm of coherence is slower for this configuration as shown in Fig. 3(g), furthur emaphasizing the intricate interplay between topology, initial conditions and the choice of stability measures.
Further analysis of Fig. 12 and Fig 13 reveals that, under both Haken–Strobl noise and intrinsic decoherence, the quantum-classical distance again identifies the complete graph and the star network with hub node initialization as the most stable configurations in all metrics. Surprisingly, despite being homogeneous, the cycle topology demonstrates the highest resilience to classicalization against the QSW decoherence. Among the remaining complex network structures, stability trends align with their degree of heterogeneity: scale-free networks exhibit higher stability than followed by small-world and Erdős–Rényi networks.
Overall, our results indicate that the star network, owing to its extreme structural heterogeneity, exhibits the highest level of quantum stability across a range of decoherence models. It is followed by the scale-free and small-world networks, which also benefit from structural heterogeneity and hub-like features. In contrast, homogeneous networks such as the cycle and Erdős–Rényi graphs, which lack prominent central nodes and have relatively low connectivity, are generally the least favorable for preserving quantum coherence, particularly under quantum stochastic noise. Notably, the complete graph, despite its homogeneity, provides high stability due to its dense connectivity, enabling sustained coherence across the network. These findings highlight the crucial role of both heterogeneity and connectivity in determining a network’s resilience to decoherence, with implications for the design of robust quantum systems.

### v2

@@ VI Conclusion
In this work, we investigate the stability of quantum systems by analyzing the dynamics of CTQWs on various network topologies under different decoherence. We define stability in terms of the preservation of quantum properties over time and examine how this preservation is influenced by the interplay between network structure and decoherence model. To analyze this, we consider both simple network structures, including star, cycle, and complete graphs, and complex topologies such as Erdős–Rényi, small-world, and scale-free networks.
@@
We evaluate the dynamics of CTQWs on these networks to examine the persistence of quantum properties, both in the absence of noise and under three decoherence models namely, Haken-Strobl, intrinsic decoherence and QSW. This systematic analysis allows us to explore how network structure and decoherence jointly influence CTQW stability. To quantify the stability, first we compute the node probability distribution of the quantum walker and the $\ell_{1}$-norm of coherence. Our simulations for node probabilities and $\ell_{1}$-norm of coherence confirm that under Haken-Strobl noise and QSW, being unital and Davies irreducible on connected networks (Appendix A), relax to the maximally mixed steady state. This contrasts with the intrinsic decoherence model, which allows for non-trivial steady states and localization. In addition to these primary measures, we also use fidelity, quantum–classical distance, and von Neumann entropy measures for providing a comprehensive characterization of the preservation of quantum properties.
@@
Based on these metrics, we find that stability varies significantly across decoherence models, intrinsic decoherence preserves coherence most effectively, followed by Haken–Strobl noise, while QSW result in the most rapid coherence loss. Our baseline scaling analysis on the star graph with initialization on center node showed that the influence of network size ($N$) is also highly decoherence model dependent: larger networks proved more coherence preservation under Haken-Strobl and intrinsic decoherence, but less stable under QSW. As expected, preservation of coherence uniformly decreased with increasing decoherence rates ($\gamma$ and $p$). Our results reveal that coherence preservation depends not only on the type of decoherence but also on the underlying topology and the specific node at which the quantum walker is initialized.
@@
Delving deeper, we observe that under Haken-Strobl noise, cycle exhibit high coherence in the absence of noise but are highly vulnerable under decoherence. In contrast, star network and complete network demonstrate greater robustness under Haken–Strobl and intrinsic decoherence, as quantified by the $\ell_{1}$-norm of coherence. In the case of QSW, star network initialized at peripheral nodes retain coherence longer and delay the onset of uniform probability distribution compared to central hub initialization. Interestingly, complete and star networks initialized at central nodes exhibit identical probability distributions and identical stability measures under Haken-Strobl and intrinsic decoherence, reflecting the impact of local connectivity due to the initialization in the fully connected hub node.
@@
For complex networks, when coherence preservation is quantified using $\ell_{1}$-norm of coherence indicates that scale-free networks are the most stable under Haken–Strobl but exhibit faster coherence decay than Erdős–Rényi networks under QSW dynamics. When stability is assessed using fidelity and von Neumann entropy, we observe distinct temporal behaviors across Erdős–Rényi, small-world, and scale-free networks. These results indicate that the presence of hubs in scale-free networks contributes to their enhanced stability. However, when using the quantum–classical distance as a stability measure, particularly under noiseless and intrinsic decoherence case, all network types yield nearly identical trajectories, making it difficult to distinguish relative stability. The interpretation of stability is inherently application dependent. For example, high entropy may impair the efficiency of quantum search algorithms [bose2000communication], while a large quantum–classical distance, indicating strong deviation from classical dynamics can benefit the design of optimized chiral quantum walks [frigerio2022quantum].
In general, our results reveal that stability is a non-trivial interplay between the network topology and the physical mechanism of the noise. High-connectivity features, such as the hubs in heterogeneous networks, provide stability against node based (Haken-Strobl) decoherence by inducing localization. However, these same features are a critical vulnerability to edge based (QSW) decoherence, where they accelerate decay of quantum coherence. Moreover, the value of coherence in noiseless scenario in such highly connected networks, whose initialization at high-degree nodes tends to promote localization but reduces coherence in noiseless conditions. This observation highlights a fundamental trade-off between localization and coherence. The implications of this trade-off are particularly relevant for quantum information processing: coherence can serve as a resource for quantum search algorithms based on quantum walks [shi2017coherence, su2018coherence], whereas high fidelity often indicative of localization can be exploited for robust quantum memories [chandrashekar2015localized]. Based on our analysis, there is significant scope for future research on stability in noisy environments by investigating additional structural properties of networks and considering the dynamics of multiple continuous-time quantum walkers (multi-CTQWs).
@@ Acknowledgments
ALJ acknowledges the University Grants Commission (UGC), India, for financial support. CM acknowledges support from the Anusandhan National Research Foundation (ANRF) India (Grants Numbers SRG/2023/001846 and EEQ/2023/001080). We are grateful to Dr. Luca Razzoli for helpful correspondence and for pointing us to his work [razzoli2022universality], which provided the analytical explanation for the universality we observed in the star initial localization on center node and complete network dynamics.
Appendix A Proof for Steady state
In this appendix we prove that the steady state of both the Haken-Strobl noise and QSW dynamics must relax to the unique maximally mixed state ($\rho_{ss}=\frac{\mathbb{I}}{N}$). This is demonstrated by verifying that their respective Liouvillians are unital and that the systems satisfy the condition for Davies irreducibility [davies1976quantum].
A.1 Steady state for Haken-Strobl Noise
From Eq. 3, Substituting the Lindblad operators which are projectors $P_{k}=|k\rangle\langle k|$ is given by:
$\frac{d\rho(t)}{dt}=\mathcal{L}(\rho)=-i[H,\rho(t)]+\gamma\sum_{k}\left(|k\rangle\langle k|\rho(t)|k\rangle\langle k|-\frac{1}{2}\{|k\rangle\langle k|k\rangle\langle k|,\rho(t)\}\right);$ (13)
For a steady state, we must have $\mathcal{L}(\rho^{ss})=0$. We check if the maximally mixed state $\rho^{ss}=c\mathbb{I}$ is a valid steady state:
$\begin{split}\mathcal{L}(c\mathbb{I})&=-i[H,c\mathbb{I}]+\gamma\sum_{k}\left(|k\rangle\langle k|\,c\mathbb{I}\,|k\rangle\langle k|-\frac{1}{2}\left\{|k\rangle\langle k|k\rangle\langle k|,\,c\mathbb{I}\right\}\right)\\ &=\gamma\sum_{k}\left(c\,|k\rangle\langle k|-\frac{1}{2}\left\{|k\rangle\langle k|,\,c\mathbb{I}\right\}\right)\\ &=\sum_{k}c|k\rangle\langle k|-c|k\rangle\langle k|=0.\end{split}$ (14)
Applying the normalization condition $Tr(\rho)=1$. We find that $Nc=1$ or $c=1/N$. Thus, the steady state density matrix is:
$\rho_{ss}=\frac{\mathbb{I}}{N}.$ (15)
This shows that $\frac{\mathbb{I}}{N}$} is a steady state solution, where N is the number of nodes.
To guarantee that this steady state is unique, we invoke the concept of Davies irreducibility [davies1976quantum]. We utilize the criteria proposed by Zhang and Barthel [zhang2024criteria], which state that a system is Davies reducible if and only if there exists a non-trivial projection $\mathcal{P}$ where (where ${\mathcal{P}}\neq 0$ and $\mathcal{P}\neq\mathbb{I}$) that satisfies two conditions: (i) $(\mathbb{I}-\mathcal{P}){P}_{k}\mathcal{P}=0$ for all Lindblad operators ${P}_{k}$. (ii) $(\mathbb{I}-\mathcal{P})(i{H}+\frac{1}{2}\sum_{k}{P}_{k}^{\dagger}{P}_{k})\mathcal{P}=0$. If we can prove that the only projections $\mathcal{P}$ satisfying these conditions are $\mathcal{P}=0$ and $\mathcal{P}=\mathbb{I}$, the quantum system is irreducible. For the first condition:
$(\mathbb{I}-\mathcal{P})|k\rangle\langle k|\mathcal{P}=0.$ (16)
This implies:
$|k\rangle\langle k|\mathcal{P}=\mathcal{P}|k\rangle\langle k|\mathcal{P}.$ (17)
Taking hermitian adjoint of this equation Eq. (17) and using the property that projectors $\mathcal{P}$ are hermitian,
$\mathcal{P}|k\rangle\langle k|={P}|k\rangle\langle k|\mathcal{P}.$ (18)
Eq. (17) and Eq. (18) imply that the commutator of $\mathcal{P}$ and $|k\rangle\langle k|$ is zero: $\mathcal{P},|k\rangle\langle k|]=0$. Evaluating the $ij^{th}$ element of the commutator for $i\neq j$:
$\langle i|[\mathcal{P}|k\rangle\langle k|]|j\rangle=\langle i|\mathcal{P}|k\rangle\langle k|j\rangle-\langle i|k\rangle\langle k|\mathcal{P}|j\rangle=\langle i|\mathcal{P}|k\rangle\delta_{kj}-\delta_{ik}\langle k|\mathcal{P}|j\rangle.$ (19)
Let us impose the $\delta_{kj}$ in the first term. Then for $i\neq j$:
$\langle i|\mathcal{P}|j\rangle\delta_{jj}-\delta_{ij}\langle j|\mathcal{P}|j\rangle=\langle i|\mathcal{P}|j\rangle-0=\mathcal{P}_{ij}=0.$ (20)
Since $\langle i|\mathcal{P}|j\rangle=0$ for all $i\neq j$, the projector $\mathcal{P}$ must be diagonal. For the second condition, substituting H=LH=L(the graph Laplacian):
$(\mathbb{I}-\mathcal{P})(i{L}+\frac{1}{2}\sum_{k}|k\rangle\langle k|k\rangle\langle k|)\mathcal{P}=(\mathbb{I}-\mathcal{P})(i{L}+\frac{1}{2}\sum_{k}|k\rangle\langle k|)\mathcal{P}=iL\mathcal{P}-iPL\mathcal{P}=0.$ (21)
This implies that $L\mathcal{P}=\mathcal{P}L\mathcal{P}$. Taking adjoint gives $\mathcal{P}L=\mathcal{P}L\mathcal{P}$. Thus, L and $\mathcal{P}$ must commute($\mathcal{P},L]$=0). Taking $ij^{th}$ element of the commutator,
$L,\mathcal{P}]_{ij}=\sum_{k}L_{ik}\mathcal{P}_{kj}-\mathcal{P}_{ik}L_{kj}=L_{ij}\mathcal{P}_{jj}-\mathcal{P}_{ii}L_{ij}=L_{ij}(\mathcal{P}_{jj}-\mathcal{P}_{ii}).$ (22)
Here we used the fact that since $\mathcal{P}$ is diagonal ($\mathcal{P}_{kj}=\mathcal{P}_{kj}\delta_{kj}$). If nodes i and j are connected, $L_{ij}=-1$. Since the commutator is zero, Eq. (22) implies that $P_{jj}=P_{ii}$ for all connected nodes. For a connected graph, this is true for all nodes, meaning all diagonal elements must be equal: $\mathcal{P}=c\mathbb{I}$.
Since the eigen values of projector can only be 0 or 1, the only possible values are $c=0$ or $c=1$. Thus $\mathcal{P}=0$ or $\mathcal{P}=\mathbb{I}$ are the only possible projections. This satisfies the condition for irreducibility. Therefore, the Haken-Strobl Lindblad operators and irreducible and the system has a unique steady state, $\rho_{ss}=\frac{\mathbb{I}}{N}$ as shown in Eq. 15. This analytical result for steady state aligns with our simulations irrespective of network topology as shown in Sec. V.1.
A.2 Steady state for Quantum Stochastic Walk
From Eq. 4, substituting the Lindblad operators for QSW $P_{kj}=|k\rangle\langle j|$. The master equation is:
$\begin{split}\frac{d\rho(t)}{dt}=\mathcal{L}(\rho)=-(1-p)i[H,\rho(t)]+p\sum_{kj}\left(L_{kj}|k\rangle\langle j|\rho(t)L_{kj}|j\rangle\langle k|-\frac{1}{2}\left\{L_{kj}|j\rangle\langle k|L_{kj}|k\rangle\langle j|\rho(t)\right\}\right).\end{split}$ (23)
Checking whether $c\mathbb{I}$ can be a valid steady state for QSW as done for Haken-Strobl noise A.1:
$\begin{split}\mathcal{L}(c\mathbb{I})&=-(1-p)i[H,c\mathbb{I}]+p\sum_{kj}\left(L^{2}_{kj}|k\rangle\langle j|\,c\mathbb{I}\,|j\rangle\langle k|-\frac{1}{2}\left\{L^{2}_{kj}|j\rangle\langle k|k\rangle\langle j|,\,c\mathbb{I}\right\}\right)\\ &=p\sum_{kj}\left(cL^{2}_{kj}|k\rangle\langle k|-\frac{1}{2}\left\{L^{2}_{kj}|j\rangle\langle j|,\,c\mathbb{I}\right\}\right).\end{split}$ (24)
where, $\frac{1}{2}\left\{L^{2}_{kj}|j\rangle\langle j|,c\mathbb{I}\right\}=cL^{2}_{kj}|j\rangle\langle j|$Substitute this in Eq. 17 gives:
$\mathcal{L}(c\mathbb{I})=0.$ (25)
Since $Tr(\rho^{ss})=1$ gives:$Nc=1$ or $c=1/N$ which means the steady state is:
$\rho_{ss}=\frac{\mathbb{I}}{N}.$ (26)
As we did in  A.1, we use the Zhang and Barthel criteria to check for irreducibility. Applying the first condition $(\mathbb{I}-\mathcal{P}){P}_{kj}\mathcal{P}=0$:
$(\mathbb{I}-\mathcal{P})L_{kj}|k\rangle\langle j|\mathcal{P}=0.$ (27)
Consider the diagonal elements of this equation. $k=j$
$(\mathbb{I}-\mathcal{P})L_{kk}|k\rangle\langle k|\mathcal{P}=0.$ (28)
This is identical to Eq. 16 for the Haken-Strobl case. It immediately implies that for QSW, the projection operator $\mathcal{P}$ must be diagonal. Now consider the off-diagonal jump operators $k\neq j$. Taking the $il^{th}$ element of this:
$\begin{split}\big((\mathbb{I}-\mathcal{P})\,L_{kj}\,|k\rangle\langle j|\,\mathcal{P}\big)_{il}&=\sum_{mn}L_{kj}\,(\mathbb{I}-\mathcal{P})_{im}\,(|k\rangle\langle j|)_{mn}\,\mathcal{P}_{nl}\\ &=\sum_{mn}L_{kj}\,(\mathbb{I}-\mathcal{P})_{im}\,\delta_{mk}\,\delta_{nj}\,\mathcal{P}_{nl}\\ &=L_{kj}\,(\mathbb{I}-\mathcal{P})_{ik}\,\mathcal{P}_{jl}\\ &=L_{kj}\,\big(1-\mathcal{P}_{kk}\big)\,\mathcal{P}_{jl}\,.\end{split}$ (29)
Here we used the result that $\mathcal{P}$ should be diagonal from Eq. 28. Also by Eq. (27), $L_{kj}(1-\mathcal{P}_{kk})\mathcal{P}_{jj}=0$. If node i and j are connected, the Laplacian elements $L_{kj}=-1$ else $L_{kj}=0$. Therefore, for connected nodes k and j: (i) If $\mathcal{P}_{jj}=1$ (node j is in the subspace), then it forces $(1-\mathcal{P}_{kk})=0$, so $\mathcal{P}_{kk}=1$ (node k is also in the same subspace). (ii) If $\mathcal{P}_{kk}=0$ (node k is out of the subspace), then $\mathcal{P}_{jj}=0$.
Since the graph is connected, either all diagonal elements are 1 ($\mathcal{P}=\mathbb{I}$) or all are 0 ($\mathcal{P}=0$). This proves that the QSW dynamics is Davies irreducible and has a unique steady state as in Eq. (26), $\rho^{ss}=\frac{\mathbb{I}}{N}$. This analytical result perfectly aligns with our numerical simulations in Sec. V.1 and  V.2, where we consistently observed convergence to this uniform distribution of probabilities with coherence decays to zero regardless of the network topology.
Appendix B Extended Stability Metrics for Simple Topologies
To obtain a more comprehensive picture of stability beyond the $\ell_{1}$-norm of coherence in Sec. V.1, we also analyze the quantum-classical distance, fidelity with the initial state, and von Neumann entropy.
Indeed, these additional metrics confirm that the complete graph and the star network (hub-initialized) consistently demonstrate the highest stability, a finding that holds for both Haken-Strobl and intrinsic decoherence. This reinforces our earlier conclusion from the $\ell_{1}$-norm of coherence analysis, which had already identified these two same networks with high degree nodes as the most robust under the Haken-Strobl model. This is evidenced by their high quantum-classical distance and fidelity, along with the lowest von Neumann entropy, as shown in Fig. 9(a),(b),(c),(e),(f),(g) and Fig. 10(a),(b). In contrast, the cycle network tends to classicalize more rapidly, indicated by its low quantum-classical distance (Fig. 9(b),(c)). However, these metrics also reveal that for the QSW model, the stability rankings invert. The cycle network, previously the least stable, now displays a slower classicalization (a higher quantum calssical distance) than the highly-connected complete and star network with center node initialization (Fig. 9(d)). Under QSW, the star network with peripheral node initialization emerges as the most stable configuration, showing the slowest decay of quantum-classical distance, high fidelity, and low entropy (Fig. 9(d),(h) and Fig. 10(c)). This is consistent with our coherence analysis, which also identified this configuration as the most robust under QSW. This emphasizes the intricate interplay between network topology, the walker’s initial condition, and the specific mechanism of decoherence, as elaborated on in the main conclusion (Sec. VI).